OpenStudy (anonymous):

Given the equation 6x squared-y=98 , find x if y = –2.

OpenStudy (anonymous):

Plug-in (replace) y with its value -2.

OpenStudy (anonymous):

(That's the first step)

OpenStudy (mertsj):

\[6x^2-y=98\] \[6x^2-(-2)=98\] \[6x^2+2=98\]

OpenStudy (mertsj):

1. Subtract 2 from both sides. 2. Divide both sides by 6 3. Take the square root of both sides and remember that each number has two square roots...a positive one and a negative one.

OpenStudy (anonymous):

After the third step, what's the answer? I'm still confused.

OpenStudy (anonymous):

There will be two answers for x. One will be positive and the other, negative.

OpenStudy (anonymous):

Show me your 3rd step.

OpenStudy (mertsj):

What did you have after you divided both sides by 6?

OpenStudy (anonymous):

6x squared = 96

OpenStudy (anonymous):

then what

OpenStudy (mertsj):

Divide both sides by 6

OpenStudy (mertsj):

1. Subtract 2 from both sides. 2. Divide both sides by 6

OpenStudy (anonymous):

6x/6= 96/6?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thats what i divide

OpenStudy (anonymous):

okay.

OpenStudy (mertsj):

\[\frac{6x^2}{6}=\frac{96}{6}\]

OpenStudy (anonymous):

in that case, 96 divided by six is 16? So, x=16?

OpenStudy (anonymous):

squared?

OpenStudy (mertsj):